OpenStudy (anonymous):

A triangle has an area of 15 in2, and two of the sides of the triangle have lengths 5 in. and 7 in. Find the angle included by these two sides. (Assume the angle is acute. Round your answer to one decimal place.)

OpenStudy (anonymous):

15 = A = (1/2)(5)(7)Sin Theta Sin Theta = (2)(15)/ (5)(7) = 30/35 = 6/7 How do I get degrees?

OpenStudy (anonymous):

in your calculator,theres somewhere \[\sin ^{-1}\] ,just type \[\sin ^{-1}\](6/7) to get your answer

OpenStudy (anonymous):

Oh... Thanks! I forgot to use the inverse.

OpenStudy (anonymous):

I got 59 degrees btw

OpenStudy (anonymous):

yes